OpenStudy (anonymous):

Please Help ... Solve using the addition method. What is the x-value of the solution to the system? 2x - 3y = -1 3x + 4y = 24 If you were to solve the system using the addition method, what would you multiply the first equation by in order to make the x's add out of this system? -2x + 3y = 6 4x - 2y = 8 When solving the system by substitution, what would you plug into the second equation for the letter x? x+2y=4 2x+2y=6 The ordered pair (4,2) is a solution to the system of inequalities below. y>3x+2 Solve the following system of equations by graphing. x + 2y = 4 2x + 4y = 8 What is the solution? Solve the following system of equations by graphing. x + 2y = 8 x + 2y = -4 What is the solution?

OpenStudy (anonymous):

multiply the top one by 4, the bottom one by 3, then add and the y terms will go bye bye

OpenStudy (dtan5457):

For clarification, addition=elimination method...right?

Nnesha (nnesha):

you can also multiply first equation by -3 and second one by 2 :D :)

OpenStudy (dtan5457):

If so.. 2x - 3y = -1 3x + 4y = 24 You would multiply the top by 4 and the bottom by 3 like satellite said. 8x-12y=-4 9x+12y=72

OpenStudy (dtan5457):

then when you add those up ^ the y is gone and you just solve for x

OpenStudy (anonymous):

Solve for x how?

OpenStudy (anonymous):

I'm so confused haha

OpenStudy (dtan5457):

8x-12y=-4 9x+12y=72 add them up^^ 17x=68

OpenStudy (dtan5457):

to make x just x, divide by 17 ^

OpenStudy (dtan5457):

so therefore x=_____?

OpenStudy (anonymous):

I got 4 ?(:

OpenStudy (dtan5457):

there you go :)

OpenStudy (anonymous):

Thanks a bunch !(: